# Constraint-Generation Policy Optimization (CGPO): Nonlinear Programming for Policy Optimization in Mixed Discrete-Continuous MDPs

### v1

@@ Introduction
We now extend the intuition above to the general deterministic MDP setting. We assume that $\pi$ can be compactly identified by a vector $\mathbf{w}\in\mathcal{W}$ of decision variables, and we shall write $V(\mathbf{w},\mathbf{s}_{1})$ to denote the value $V(\pi_{\mathbf{w}},\mathbf{s}_{1})$ of the policy $\pi_{\mathbf{w}}$ parameterized by $\mathbf{w}$. In what follows, we focus our attention on approximate policy sets $\tilde{\Pi}_{\mathcal{W}}=\{\pi_{\mathbf{w}}:\mathbf{w}\in\mathcal{W}\}$ that can be compactly described by piecewise functions (1).
@@
First, observe that for every possible initial state $\mathbf{s}_{1}$, there exists a fixed optimal plan $\alpha^{*}=[\mathbf{a}_{1}^{*},\dots\mathbf{a}_{T}^{*}]$ such that $V(\alpha^{*},\mathbf{s}_{1})=V(\pi_{H}^{*},\mathbf{s}_{1})$. On the other hand, since we only have access to an expressivity class of approximate policies $\tilde{\Pi}_{\mathcal{W}}$, the error $\varepsilon(\mathbf{w},\mathbf{s}_{1})$ of $\pi_{\mathbf{w}}$ in state $\mathbf{s}_{1}$ relative to $V(\pi_{H}^{*},\mathbf{s}_{1})$ (according to (2)) must be
$\varepsilon(\mathbf{w},\mathbf{s}_{1})\geq\max_{\alpha\in\mathcal{A}^{T}}V(\alpha,\mathbf{s}_{1})-V(\mathbf{w},\mathbf{s}_{1}),$ (3)
and thus the worst-case error $\varepsilon(\mathbf{w})$ of $\mathbf{w}$ is
@@
However, since we seek the approximate optimal policy $\tilde{\pi}^{*}\in\tilde{\Pi}_{\mathcal{W}}$, we can directly minimize (4) over $\mathcal{W}$, obtaining the following infinitely-constrained mixed-integer program:
$\min_{\mathbf{w}\in\mathcal{W},\,\varepsilon\in[0,\infty)}\quad$ $\varepsilon$ (5)
@@
However, the constraint is highly nonlinear, turning (5) into a bi-level program and making analysis of this particular formulation difficult. Instead, since the $\max$’s must hold for all states and actions, we can rewrite the problem (5) as:
$\min_{\mathbf{w}\in\mathcal{W},\,\varepsilon\in[0,\infty)}\quad$ $\varepsilon$ (6)
s.t. $\varepsilon\geq V(\alpha,\mathbf{s}_{1})-V(\pi,\mathbf{s}_{1})$
$\quad\,\forall\alpha\in\mathcal{A}^{T},\,\mathbf{s}_{1}\in\mathcal{S}_{1}.$
The goal is therefore to solve (6), which is still infinitely-constrained when $\mathcal{S}$ or $\mathcal{A}$ are infinite spaces. Instead, we solve it by splitting the $\min$ over $\mathcal{W}$ and the $\max$ over $(\alpha,\mathbf{s}_{1})$ into two problems, and apply constraint generation (Blankenship and Falk 1976; Chembu et al. 2023). Specifically, starting with a fixed arbitrary scenario $(\mathbf{s}_{1},\alpha)$, we form the constraint set $\mathcal{C}=\{(\mathbf{s}_{1},\alpha)\}$, and then repeatedly solve the following two problems:
  • Outer Problem: Solve (6) over the finite constraint set $\mathcal{C}$ to obtain policy $\mathbf{w}_{1}^{*}\in\mathcal{W}$.
  • Inner Problem: Solve (4):
    $(\mathbf{s}_{1}^{*},\alpha^{*})\in{\arg\!\max}_{\mathbf{s}_{1}\in\mathcal{S}_{1},\alpha\in\mathcal{A}^{T}}\left[V(\alpha,\mathbf{s}_{1})-V(\mathbf{w}_{0}^{*},\mathbf{s}_{1})\right],$
    for the highest-error scenario for $\mathbf{w}_{1}^{*}$, and append it to $\mathcal{C}$.
These two steps are repeated until the constraint added to the outer problem is no longer binding, i.e. $V(\alpha^{*},\mathbf{s}_{1}^{*})-V(\mathbf{w}_{t}^{*},\mathbf{s}_{1}^{*})\leq 0$, since the solution of the outer problem will not change with the addition of the constraint. We name this approach Constraint-Generation Policy Optimization (CGPO) (see Figure 1). Intuitively, we can view CGPO as a policy iteration algorithm where the inner problem adversarially critiques the policy with a worst-case trajectory and the outer problem improves the policy w.r.t. all critiques.
@@
Upon termination, GCPO’s use of constraint generation guarantees an optimal (i.e., lowest error) policy within the specified policy class (see Theorem 1). Further, while we cannot provide a general finite-time guarantee of termination with an optimal policy, CGPO is an anytime algorithm that provides a best policy and a bound on performance at each iteration. Remarkably, CGPO is applicable for solving infinite state/action bi-level policy optimization problems by exploiting the fact that only several constraints may be active at the optimal solution. In addition, at each stage, the solution to the inner problem allows us to interpret and analyze the state with the worst-case policy performance with respect to $\pi$ (cf. Figure 5), and generate a counterfactual explanation of what actions should have been made.
Chance Constraints for Stochastic MDPs
We now turn our attention to MDPs with stochastic transitions, which are considerably more difficult to handle since they cannot be reformulated directly in the form (5).
A number of approaches could be used to encode stochastic variables, including sample approximation (SAA) (Kleywegt, Shapiro, and Homem-de Mello 2002) and hindsight encoding (Raghavan et al. 2017) on the inner constraints, which would typically lead to a blowup in the problem size. Chance-constrained approaches (Ono et al. 2015; Farina, Giulioni, and Scattolini 2016; Ariu et al. 2017) instead generate constraints that hold with high probability, which often alleviate the computational burden at the expense of introducing some degree of (controllable) bias. Motivated by the recent applications of chance-constraints in planning and control, we take the latter approach, which will lead to a particularly elegant bi-level reformulation of (6) as we now see.
To achieve this, we will require that the state transition function $P(\mathbf{s},\mathbf{a},\mathbf{s}^{\prime})$ has a natural reparameterization as a deterministic function of $(\mathbf{s}$, $\mathbf{a})$ of some exogenous i.i.d. noise variable $\xi$ with density $q(\cdot)$ on support $\Xi$, e.g. $\mathbf{s}^{\prime}=g(\mathbf{s},\mathbf{a},\xi)$ (Bueno et al. 2019). Given a “confidence” threshold $p\in(0,1)$ close to 1, we also assume the existence of a computable interval $\Xi_{p}=[\xi_{l},\xi_{u}]$ such that $\mathbb{P}\!\left(\xi_{l}\leq\xi\leq\xi_{u}\right)\geq p$.
Thus, we can repeat the reasoning of the previous section, by considering the worst case not only over $\mathbf{s}_{1}$, but also over possible trajectories or rollouts $\xi_{1:T}=[\xi_{1},\dots\xi_{T}]\in\Xi_{p}^{T}$. The analogue of (5) in the stochastic setting is:
$\min_{\mathbf{w}\in\mathcal{W},\,\varepsilon\in[0,\infty)}\quad$ $\varepsilon$ (7)
s.t. $\varepsilon\geq\max_{\mathbf{s}_{1}\in\mathcal{S}_{1}}\max_{\xi_{1:T}\in\Xi_{p}^{T}}\max_{\alpha\in\mathcal{A}^{T}}\big{[}$
$\quad V(\alpha,\mathbf{s}_{1},\xi_{1:T})-V(\mathbf{w},\mathbf{s}_{1},\xi_{1:T})\big{]},$
where $V(\cdot,\mathbf{s}_{1},\xi_{1:T})$ corresponds to the total reward of the policy or plan accumulated over trajectory $\xi_{1:T}$ starting in $\mathbf{s}_{1}$. In the stochastic setting, it is important to note that $\varepsilon$ only holds with probability $p^{T}$ for a horizon $T$ problem. In practice, we choose $p$ such that $p^{T}=\alpha$ equates to some desired global probability bound on the entire planning trajectory $\alpha$.
@@
$\min_{\mathbf{w}\in\mathcal{W},\,\varepsilon\in[0,\infty)}\quad$ $\varepsilon$ (8)
s.t. $\varepsilon\geq V(\alpha,\mathbf{s}_{1},\xi_{1:T})-V(\mathbf{w},\mathbf{s}_{1},\xi_{1:T}),$
$\quad\,\forall\alpha\in\mathcal{A}^{T},\,\mathbf{s}_{1}\in\mathcal{S}_{1},\,\xi_{1:T}\in\Xi_{p}^{T}.$
Therefore, we can again apply constraint generation to solve this problem in two stages, in which the inner optimization produces not only a worst-case initial state and action sequence, but also the disturbances $\xi_{1:T}^{*}$ that reproduce all future worst-case state realizations $\mathbf{s}_{2},\dots\mathbf{s}_{T}$. A complete description of this algorithm is provided in Algorithm 1.
@@
  Initialize $p\in(0,1)$ and compute $\Xi_{p}$
  Initialize $\mathbf{s}_{1}\in\mathcal{S}_{1}$, $\xi_{1:T}\in\Xi_{p}^{T}$ and $\alpha\in\mathcal{A}^{T}$ arbitrarily
  Initialize constraint set $\mathcal{C}_{0}=\{(\mathbf{s}_{1},\xi_{1:T},\alpha)\}$
  Set $t=0$
@@
$\mathrm{s.t.}\quad$ $\varepsilon\geq V(\alpha,\mathbf{s}_{1},\xi_{1:T})-V(\mathbf{w},\mathbf{s}_{1}),$
$\forall\,(\mathbf{s}_{1},\xi_{1:T},\alpha)\in\mathcal{C}_{t}$
  Step 1: Solve the inner problem:
$(\mathbf{s}_{1}^{*},\xi_{1:T}^{*},\alpha^{*})\in{\arg\!\max}_{\mathbf{s}_{1}\in\mathcal{S}_{1},\,\xi_{1:T}\in\Xi_{p}^{T},\,\alpha\in\mathcal{A}^{T}}\big{[}$
$\qquad V(\alpha,\mathbf{s}_{1},\xi_{1:T})-V(\mathbf{w}_{t}^{*},\mathbf{s}_{1},\xi_{1:T})\big{]}$
@@
  if $V(\alpha^{*},\mathbf{s}_{1}^{*},\xi_{1:T}^{*})-V(\mathbf{w}_{t}^{*},\mathbf{s}_{1}^{*},\xi_{1:T}^{*})\leq 0$ then
     Terminate with policy $\mathbf{w}_{t}^{*}$ and error $\varepsilon_{t}^{*}$
@@
     Set $\mathcal{C}_{t+1}=\mathcal{C}_{t}\cup\{(\mathbf{s}_{1}^{*},\xi_{1:T}^{*},\alpha^{*})\}$
     Set $t=t+1$ and go to Step 1
@@
Remarks
In discrete-state problems, we can instead analytically marginalize out the exogenous noise to produce an “expected” transition model $\mathbf{s}^{\prime}=\sum_{\xi\in\Xi}g(\mathbf{s},\mathbf{a},\xi)q({\xi})$. For distributions on an infinite support, i.e. $\mathbb{Z}$ or $\mathbb{N}$, it is possible to marginalize over a truncated noise distribution $\tilde{q}({\xi})\propto q(\xi)1\{\xi\leq\xi_{p}\}$, where $p\in(0,1)$ is a parameter close to 1, and $\xi_{p}$ is chosen such that $\mathbb{P}\!\left(\xi\leq\xi_{p}\right)\geq p$.
Discussion of Convergence
Finite-time termination of Algorithm 1 is impossible to show without strong assumptions (i.e. convexity). However, we can still show under mild assumptions that, if Algorithm 1 terminates at some iteration $t$, then $\mathbf{w}_{t}^{*}$ is optimal in $\mathcal{W}$ (with probability $p^{T}$ in the stochastic case).
@@
If $\mathcal{S}_{1}$, $\mathcal{A}$, $\Xi_{p}$ and $\mathcal{W}$ are non-empty compact subsets of Euclidean space, $V(\alpha,\mathbf{s},\xi_{1:T})$ and $V(\mathbf{w},\mathbf{s},\xi_{1:T})$ are continuous, and Algorithm 1 terminates at iteration $t$, then $\mathbf{w}_{t}^{*}$ is optimal for problem (8).
@@
The proof is adapted from Theorem 2.1 in Blankenship and Falk (1976), as discussed in the Appendix. ∎
Remarks
Termination of CGPO guarantees that $\mathbf{w}_{t}$ is optimal in $\tilde{\Pi}_{\mathcal{W}}$, and $\varepsilon_{t}^{*}$ is the corresponding optimality gap. Moreover, if $\varepsilon_{t}^{*}=0$ and the MDP is deterministic, then $\pi_{\mathbf{w}^{*}}$ is the optimal policy for the MDP. This means we can use the gap estimate as a principled way to compare and validate different policy classes in practice. Blankenship and Falk (1976) also provides a modified approach (Algorithm 2.2) identical to CGPO, except for the removal of constraints that become non-binding from the outer problem at each iteration. If the problem $\max_{\mathbf{w}\in\mathcal{W}}V(\mathbf{w},\mathbf{s}_{1})$ is convex, convergence of CGPO could be established.
Analysis of Problem Expressivity Classes
In Table 1, we discuss the relationship between the expressivity classes of policies and state transition dynamics and the resulting classes of the inner and outer optimization problems. The policy classes of interest include PWS-C, PWL-C, PWS-L, PWL-L, PWN-N and the different variants of piecewise polynomial policies. Meanwhile, expressivity classes for state dynamics and reward include linear, polynomial and general nonlinear functions (their piecewise counterparts generally fall under the same expressivity classes, and are excluded for brevity).
Interestingly, when the policy and state dynamics are both linear, the outer problem is a PP. To see this, consider the transition function $s^{\prime}=s+a$, and the policy $a=ws$. Then:
$s^{\prime}$ $=s+a=s+ws=(w+1)s$
$a^{\prime}$ $=ws^{\prime}=w(w+1)s=O(w^{2})$
$s^{\prime\prime}$ $=s^{\prime}+a^{\prime}=(w+1)s+w(w+1)s=(w+1)^{2}s$
$a^{\prime\prime}$ $=ws^{\prime\prime}=w(w+1)^{2}s=O(w^{3}),$
which are polynomial in $w$. In a similar vein, a PWL-C policy and linear dynamics result in a MIBCP outer problem due to the bilinear interaction between successor state decision variables and policy weights in the linear conditions.
Our experiments in the next section empirically evaluate PWS-C and PWS-L policies with linear dynamics, and Q policies with nonlinear dynamics. This requires solving mixed-integer problems with large numbers of decision variables ranging from MILP to PP to MINLP. However, as algorithmic improvements in nonlinear optimizers continue to improve their search efficiency, we expect the performance of CGPO will also improve accordingly, allowing solutions for larger problems and policy expressivity classes.

### v2

@@ 3.1 Constraint Generation for Deterministic DC-MDPs
We assume $\pi$ can be compactly identified by a vector $\mathbf{w}\in\mathcal{W}$ of decision variables, and we use the shorthand $V(\mathbf{w},\mathbf{s}_{1})$ to denote the value $V(\pi_{\mathbf{w}},\mathbf{s}_{1})$ of policy $\pi_{\mathbf{w}}$ parameterized by $\mathbf{w}$. We focus our attention on approximate policy sets $\tilde{\Pi}_{\mathcal{W}}=\{\pi_{\mathbf{w}}:\mathbf{w}\in\mathcal{W}\}$ enumerated by a compact set of parameters $\mathcal{W}$.
@@
First, observe that for every possible initial state $\mathbf{s}_{1}$, there exists a fixed optimal plan $\alpha^{*}=[\mathbf{a}_{1}^{*},\dots\mathbf{a}_{T}^{*}]$ such that $V(\alpha^{*},\mathbf{s}_{1})=V(\pi_{H}^{*},\mathbf{s}_{1})$. On the other hand, since we only have access to an expressivity class of approximate policies $\tilde{\Pi}_{\mathcal{W}}$, the error $\varepsilon(\mathbf{w},\mathbf{s}_{1})$ of $\pi_{\mathbf{w}}$ in state $\mathbf{s}_{1}$ relative to $V(\pi_{H}^{*},\mathbf{s}_{1})$ according to (1) must be
$\varepsilon(\mathbf{w},\mathbf{s}_{1})\geq\max_{\alpha\in\mathcal{A}^{T}}V(\alpha,\mathbf{s}_{1})-V(\mathbf{w},\mathbf{s}_{1}),$
and thus the (global) worst-case error $\varepsilon(\mathbf{w})$ of $\mathbf{w}$ is
@@
However, since we seek the approximate optimal policy $\tilde{\pi}^{*}\in\tilde{\Pi}_{\mathcal{W}}$, we can directly minimize (2) over $\mathcal{W}$, obtaining the infinitely-constrained mixed-integer program:
$\min_{\mathbf{w}\in\mathcal{W},\,\varepsilon\in[0,\infty)}$ $\varepsilon$ (3)
@@
However, the constraint is highly nonlinear, turning (3) into a bilevel program and making analysis of this particular formulation difficult. Instead, since the $\max$’s must hold for all states and actions, we can rewrite the problem (3) as:
$\min_{\mathbf{w}\in\mathcal{W},\,\varepsilon\in[0,\infty)}$ $\varepsilon$ (4)
s.t. $\varepsilon\geq V(\alpha,\mathbf{s}_{1})-V(\pi,\mathbf{s}_{1}),\quad\,\forall\alpha\in\mathcal{A}^{T},\,\mathbf{s}_{1}\in\mathcal{S}_{1}.$
The goal is therefore to solve (4), which is still infinitely-constrained when $\mathcal{S}$ or $\mathcal{A}$ are infinite spaces. Instead, we solve it by splitting the $\min$ over $\mathcal{W}$ and the $\max$ over $(\alpha,\mathbf{s}_{1})$ into two problems, and apply constraint generation (Blankenship and Falk, 1976; Chembu et al., 2023).
Specifically, starting with a fixed arbitrary scenario $(\mathbf{s}_{1},\alpha)$, we form the constraint set $\mathcal{C}=\{(\mathbf{s}_{1},\alpha)\}$, and then solve the following two problems:
Outer
Solve (4) within the set of finite constraints $\mathcal{C}$ to obtain $\mathbf{w}^{*}\in\mathcal{W}$.
Inner
Solve (2) ${\arg\!\max}_{\mathbf{s}_{1}\in\mathcal{S}_{1},\alpha\in\mathcal{A}^{T}}\!\left[V(\alpha,\mathbf{s}_{1})\!-\!V(\mathbf{w}^{*},\mathbf{s}_{1})\right]$, for the highest-error scenario for $\mathbf{w}^{*}$, and append it to $\mathcal{C}$.
These two steps are repeated until the constraint added to the outer problem is no longer binding, i.e. $V(\alpha^{*},\mathbf{s}_{1}^{*})-V(\mathbf{w}^{*},\mathbf{s}_{1}^{*})\leq 0$, since the solution of the outer problem will not change with the addition of the new constraint. We name our approach Constraint-Generation Policy Optimization (CGPO). In retrospect, we can view CGPO as a policy iteration algorithm where the inner problem adversarially critiques the policy with a worst-case trajectory and the outer problem improves the policy w.r.t. all critiques.
@@
Upon termination, CGPO guarantees an optimal (lowest error) policy within the specified policy class (Theorem 1). While we cannot provide a general finite-time guarantee of termination with an optimal policy, CGPO is an anytime algorithm that provides a best policy and a bound on performance at each iteration. At each iteration, the solution to the inner problem allows analysis of the worst-case trajectory with respect to $\pi$ (Fig. 7), and generates a counterfactual explanation of what actions should have been made.
3.2 Chance Constraints for Stochastic DC-MDPs
Chance-constrained optimization (Ariu et al., 2017; Farina et al., 2016; Ono et al., 2015) allows us to derive high probability intervals on stochastic MDP transitions and reduce our solution to a robust optimization problem with probabilistic performance guarantees. To achieve this, we will require that the state transition function $P(\mathbf{s},\mathbf{a},\mathbf{s}^{\prime})$ has a natural reparameterization as a deterministic function $g$ of $(\mathbf{s}$, $\mathbf{a})$ and some exogenous independent and identically distributed noise variable $\xi$ with density $q(\cdot)$ on support $\Xi$, e.g. $\mathbf{s}^{\prime}=g(\mathbf{s},\mathbf{a},\xi)$ (Bueno et al., 2019; Patton et al., 2022). Given a threshold $p\in(0,1)$ close to 1, we also assume the existence of a computable interval $\Xi_{p}=[\xi_{l},\xi_{u}]$ such that $\mathbb{P}\!\left(\xi_{l}\leq\xi\leq\xi_{u}\right)\geq p$.
Thus, we can repeat the derivation of (3) by considering the worst case not only over $\mathbf{s}_{1}$, but also over possible noise variables $\xi_{1:T}=[\xi_{1},\dots\xi_{T}]\in\Xi_{p}^{T}$. The final problem is:
$\min_{\mathbf{w}\in\mathcal{W},\,\varepsilon\in[0,\infty)}$ $\varepsilon$ (5)
s.t. $\varepsilon\!\geq\!\max_{\mathbf{s}_{1}\in\mathcal{S}_{1}}\max_{\xi_{1:T}\in\Xi_{p}^{T}}\max_{\alpha\in\mathcal{A}^{T}}\left[V(\alpha,\mathbf{s}_{1},\xi_{1:T})\!-\!V(\mathbf{w},\mathbf{s}_{1},\xi_{1:T})\right],$
where $V(\cdot,\mathbf{s}_{1},\xi_{1:T})$ corresponds to the total reward of the policy or plan accumulated over trajectory $\xi_{1:T}$ starting in $\mathbf{s}_{1}$. In the stochastic setting, $\varepsilon$ only holds with probability $p^{T}$ for a horizon $T$ problem. Experimentally, we found that choosing $p$ such that $p^{T}=P$, where $P$ is a desired probability bound on the full planning trajectory, was sufficient [^1] Hence, we arbitrarily chose $p=0.995$ for our empirical evaluations..
@@ 3.2 Chance Constraints for Stochastic DC-MDPs
$\min_{\mathbf{w}\in\mathcal{W},\,\varepsilon\in[0,\infty)}$ $\varepsilon$ (6)
s.t. $\varepsilon\geq V(\alpha,\mathbf{s}_{1},\xi_{1:T})-V(\mathbf{w},\mathbf{s}_{1},\xi_{1:T}),\quad\,\forall\alpha\in\mathcal{A}^{T},\,\mathbf{s}_{1}\in\mathcal{S}_{1},\,\xi_{1:T}\in\Xi_{p}^{T}.$
Therefore, we can again apply constraint generation to solve this problem in two stages, in which the inner optimization produces not only a worst-case initial state and action sequence, but also the disturbances $\xi_{1:T}^{*}$ that reproduce all future worst-case state realizations $\mathbf{s}_{2},\dots\mathbf{s}_{T}$. The overall workflow of the computations is summarized as Algorithm 1.
@@
  Initialize $p\in(0,1)$, $\Xi_{p}$, $\mathbf{s}_{1}\in\mathcal{S}_{1}$, $\xi_{1:T}\in\Xi_{p}^{T}$, $\alpha\in\mathcal{A}^{T}$
  Set $t=0$ and $\mathcal{C}_{0}=\{(\mathbf{s}_{1},\xi_{1:T},\alpha)\}$
@@
$\mathrm{s.t.}$ $\varepsilon\geq V(\alpha,\mathbf{s}_{1},\xi_{1:T})-V(\mathbf{w},\mathbf{s}_{1}),\quad\,\forall\,(\mathbf{s}_{1},\xi_{1:T},\alpha)\in\mathcal{C}_{t}$
  Step 2: Solve the inner problem:
$(\mathbf{s}_{1}^{*},\xi_{1:T}^{*},\alpha^{*})\in{\arg\!\max}_{\mathbf{s}_{1}\in\mathcal{S}_{1},\,\xi_{1:T}\in\Xi_{p}^{T},\,\alpha\in\mathcal{A}^{T}}\left[V(\alpha,\mathbf{s}_{1},\xi_{1:T})-V(\mathbf{w}_{t}^{*},\mathbf{s}_{1},\xi_{1:T})\right]$
@@
  if $V(\mathbf{w}_{t}^{*},\mathbf{s}_{1}^{*},\xi_{1:T}^{*})\geq V(\alpha^{*},\mathbf{s}_{1}^{*},\xi_{1:T}^{*})$ (or $\|\mathbf{w}_{t}^{*}-\mathbf{w}_{t-1}^{*}\|<\delta$then
     Terminate with policy $\pi_{\mathbf{w}_{t}^{*}}$ and error $\varepsilon_{t}^{*}$
@@
     Set $t=t+1$, $\mathcal{C}_{t+1}=\mathcal{C}_{t}\cup\{(\mathbf{s}_{1}^{*},\xi_{1:T}^{*},\alpha^{*})\}$ and go to Step 1
@@
Remark
Algorithm 1 may not terminate in general, unless convexity assumptions are placed on $V(\alpha,\mathbf{s})$ or $V(\mathbf{w},\mathbf{s})$. Thus, in typical applications, one would terminate when $\|\mathbf{w}_{t}^{*}-\mathbf{w}_{t-1}^{*}\|<\delta$ for chosen hyper-parameter $\delta$. However, our experiments have shown empirically that Algorithm 1 terminates with an exact optimal solution for most policy classes considered, at which point both the return and error curves “plateau” (cf. Fig. 3).
3.3 Convergence
Under mild conditions, if CGPO terminates at some iteration $t$, then $\mathbf{w}_{t}^{*}$ is optimal in $\mathcal{W}$ (with probability at least $p^{T}$ in the stochastic case).
@@ 3.3 Convergence
If $\mathcal{S}_{1}$, $\mathcal{A}$, $\Xi_{p}$ and $\mathcal{W}$ are non-empty compact subsets of Euclidean space, $V(\alpha,\mathbf{s},\xi_{1:T})$ and $V(\mathbf{w},\mathbf{s},\xi_{1:T})$ are continuous, and CGPO terminates at iteration $t$, then $\mathbf{w}_{t}^{*}$ is optimal for problem (6).
@@
In the notation of (Blankenship and Falk, 1976), we define $X^{0}=[0,2BT]$, $X=\mathcal{W}$, $Y=\mathcal{A}^{T}\times\mathcal{S}_{1}\times\Xi_{p}^{T}$. Making the change of variable $\varepsilon=x_{0}\in X^{0}$, $x=\mathbf{w}$, and $y=(\alpha,\mathbf{s}_{1},\xi_{1:T})\in Y$, and defining the continuous function $f(x,y)=\varepsilon(\alpha,\mathbf{s}_{1},\xi_{1:T})=V(\alpha,\mathbf{s}_{1},\xi_{1:T})-V(\mathbf{w},\mathbf{s}_{1},\xi_{1:T})$, the problem (6) is equivalent to the problem:
$\min_{(x_{0},x)\in X^{0}\times X}\,x_{0}\qquad\mathrm{s.t.}\quad f(x,y)-x_{0}\leq 0,\quad\forall y\in Y$
as discussed on p. 262 of (Blankenship and Falk, 1976). Similarly, Algorithm 1 can be reparameterized as Algorithm 2.1 in the aforementioned paper, with $x_{t}=\mathbf{w}_{t}^{*}$. Thus, the assumption of Theorem 2.1 in (Blankenship and Falk, 1976) holds and $\mathbf{w}_{t}^{*}$ is a solution of (6) as claimed. ∎
Remark
Termination of CGPO guarantees that $\mathbf{w}_{t}^{*}$ is optimal in $\tilde{\Pi}_{\mathcal{W}}$, and $\varepsilon_{t}^{*}$ is the corresponding optimality gap. Moreover, if $\varepsilon_{t}^{*}=0$ and the MDP is deterministic, then $\pi_{\mathbf{w}_{t}^{*}}$ is the optimal policy in $\Pi$. This suggests the gap estimate can be used as a principled way to assess the suitability of different policy classes.
3.4 Problem Expressivity Class Analysis
Dynamics / Reward